A positive charge Q, located at the origin, produces an electric potential V = kQ/r at point r. A point charge q is moved from 4r to r. The increase in its electric potential energy is

Question 19 options:

a)

2kQq/(3r).

b)

kQq/(2r).

c)

kQq/(4r).

d)

kQq/(3r).

e)

kQq/r.

f)

3kQq/(4r).
Answer #1

∆U = kQq( 1/r2 - 1/r1)

=) ∆U = kQq ( 1/r - 1/4r)

=) ∆U = kQq ( 3r/4 ) = 3kQq/4r.

Option f) is correct .